Fermi National Laboratory
 

Accelerator Update for June 20, 2006 - June 21, 2006

 

Tuesday June 20

The day shift began with Operations monitoring store 4781 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) not stacking due to a LCW problem with HV102, with the Booster, Main Injector (MI), and Recycler conducting studies, with MiniBooNE off due to a Horn power supply ground fault, and with NuMI taking beam.

At 9:20 AM, a Linac expert reported that a quadrupole magnet (QPS511) was shorted. This held off all beam except for the Neutron Therapy Facility (NTF). Technicians will install an isolation block and he will tune around the magnet when beam returns.

Operations prepared Linac for access at 10:27 AM.

Operations terminated store 4781 at 10:40 AM.

A MiniBooNE expert reported at 10:42 AM that techs had repaired the Horn power supply and they were ready for beam.

The Linac access ended and beam returned at 11:08 AM.

Operations put the TeV into shot setup at 11:13 AM.

Operations established store 4782 at 2 PM with an initial luminosity of 111.37E30.

A Linac expert tuned up the beam line at 2:32 PM.

A MI expert resumed a study period at 3:53 PM.

At 3:54 PM, Switchyard (SY) experts began trying to bring the beam line up, but problems were hindering them. A power supply expert investigated and discovered he needed an access to connect a cable; this will occur on Wednesday.

At 5:09 PM, Operations reported that a quadrupole power supply problem (I:Q712) was also holding SY off. Experts investigated.

Operations began preparing the Recycler at 7:37 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 8:31 PM.

At 11:59 PM, the TeV aborted due to a sector C3 vacuum indication, but operators could not find a vacuum problem. There was no quench. Operations investigated.

Wednesday June 21

The midnight shift began with Operations monitoring the stash, with Pbar stacking, with Operations investigating the C3 abort, and with MiniBooNE and NuMI taking beam.

Operations concluded at 1:03 AM that it was probably a controls problem. Operations called in a controls expert.

Operations began preparing the Recycler at to stash at 1:25 AM. The stash ended and Pbar resumed stacking around 2:47 AM.

The controls expert reported at 3:39 AM that he replaced a C200 module.

A TeV expert began a study period at 4:15 AM.

Operation prepared D0 for an access at 6:25 AM.

Operations turned off NuMI at 6:31 AM for an access later in the morning.

The Recycler began preparing to stash at 7:22 AM.

The Plans for Wednesday

The plans for today are to stack and stash. D0 has begun its 16-hour access. As soon as D0 completes its access, Operations will establish a store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/21/2006    email Fermilab